# On the stability and conditioning of a fictitious domain formulation for fluid-structure interaction problems

### v1

@@ 9. Numerical tests
Since the definition of ${\bm{\Lambda}}$ changes with the choice of coupling term $\mathbf{c}$, we point out that if $\mathbf{c}=\mathbf{c}_{1}$ then the error computation for $\bm{\lambda}$ is done in the standard $\mathbf{H}^{1}(\mathcal{B})$ norm. On the other hand, when $\mathbf{c}=\mathbf{c}_{0}$ the error is computed by using the norm of the dual space $(\mathbf{H}^{1}(\mathcal{B}))^{\prime}$: such norm is evaluated by solving the auxiliary problem $-\Delta\bm{\psi}+\bm{\psi}=\bm{\lambda}-\bm{\lambda}_{h}$ with homogeneous Neumann boundary conditions and by computing the $H^{1}$-norm of $\bm{\psi}$.
@@ 9.1. The shifted square
$\bm{\lambda}(s_{1},s_{2})=(e^{s_{1}},e^{s_{2}})$ for $\qquad\text{for }\mathbf{s}=(s_{1},s_{2})\in\mathcal{B},$
@@
(45) $\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-4}$ if $\quad\text{if }\mathbf{c}=\mathbf{c}_{0}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{0,h},$
$\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-2}$ if $\quad\text{if }\mathbf{c}=\mathbf{c}_{1}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{1,h}.$
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 9.2.1. The disk
$p(x,y)=\begin{cases}\sin(\pi x)\sin(\pi y)-{4}/{\pi^{2}}-|\Omega^{s}|/(2|\Omega^{f}|)&\text{in }\Omega^{f}\\ \sin(\pi x)\sin(\pi y)-{4}/{\pi^{2}}+1/2&\text{in }\Omega^{s}=\mathcal{B}\end{cases}$
@@ 9.2.3. The stretched annulus
=$\mathcal{B}=\left\{\mathbf{s}\in\mathbb{R}^{2}:\frac{1}{16}\leq\left(s_{1}-\frac{1}{2}\right)^{2}+\left(s_{2}-\frac{1}{2}\right)^{2}\leq\frac{1}{64}\right\},$
@@
Figure 10. Condition number of disk, flower, and annulus tests.
@@
Concerning the disk example, we observe that $p$ is discontinuous along the interface, thus ${p\in H^{s}(\Omega)}$ with $20\leq s<1/$ 2. Hence, we cannot expect optimal rate of convergence for this variable even if the coupling term is computed exactly. In the other two cases, since the solution is regular, the method shows optimal convergence. The only critical case is given by the error estimate of the variable $\lambda$ in Figure 13 that has a surprising behavior in the last level of refinement. This might be due to the ill-conditioning of the matrix: according to the proven estimate, the condition number is growing like $O(h^{-4})$ and in the last test we might have reached the critical value of $h$ so that the error stops decreasing.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11. Error convergence for the disk test.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12. Error convergence for the flower-shaped solid test.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption

### v2

@@ 6. Numerical tests
Since the definition of $\bm{\Lambda}$} changes with the choice of coupling term $\mathbf{c}$, we point out that if $\mathbf{c}=\mathbf{c}_{1}$ then the error computation for $\bm{\lambda}$ is done in the standard $\mathbf{H}^{1}(\mathcal{B})$ norm. On the other hand, when $\mathbf{c}=\mathbf{c}_{0}$ the error is computed by using the norm of the dual space $(\mathbf{H}^{1}(\mathcal{B}))^{\prime}$: such norm is evaluated by solving the auxiliary problem $\Delta\bm{\psi}+\bm{\psi}=\bm{\lambda}-\bm{\lambda}_{h}$ with homogeneous Neumann boundary conditions and by computing the $H^{1}$-norm of $\bm{\psi}$.
@@ 6.1. The shifted square
$\bm{\lambda}(s_{1},s_{2})=(e^{s_{1}},e^{s_{2}})$ $\qquad\text{for }\mathbf{s}=(s_{1},s_{2})\in\mathcal{B},$
@@
(45) $\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-4}$ $\quad\text{if }\mathbf{c}=\mathbf{c}_{0}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{0,h},$
$\mathsf{cond}(\mathcal{A}_{\square})\leq Ch^{-2}$ $\quad\text{if }\mathbf{c}=\mathbf{c}_{1}\text{ or }\mathbf{c}_{h}=\mathbf{c}_{1,h}.$
@@ 6.2. Disk, flower and stretched annulus
Figure 10. Condition number of disk, flower, and annulus tests.
@@ 6.2.1. The disk
[…]
@@ 6.2.3. The stretched annulus
$\mathcal{B}=\left\{\mathbf{s}\in\mathbb{R}^{2}:\frac{1}{16}\leq\left(s_{1}-\frac{1}{2}\right)^{2}+\left(s_{2}-\frac{1}{2}\right)^{2}\leq\frac{1}{64}\right\},$
@@
Figure 12. Error convergence for the flower-shaped solid test.
@@
Concerning the disk example, we observe that $p$ is discontinuous along the interface, thus $p\in H^{s}(\Omega)$} with $0\leq s<1/2$. Hence, we cannot expect optimal rate of convergence for this variable even if the coupling term is computed exactly. In the other two cases, since the solution is regular, the method shows optimal convergence. The only critical case is given by the error estimate of the variable $\lambda$ in Figure 13 that has a surprising behavior in the last level of refinement. This might be due to the ill-conditioning of the matrix: according to the proven estimate, the condition number is growing like $O(h^{-4})$ and in the last test we might have reached the critical value of $h$ so that the error stops decreasing.
6.3. Time dependent problem: the stretched annulus
In this last subsection, we report some numerical results for the time dependent Problem 1 of a stretched annulus immersed in a square fluid domain. We focus on the effect of small cut cells on the stability in time.
Let $\Omega$ be the unit square and $\mathcal{B}=\{\mathbf{s}\in\mathbb{R}^{2}:\,0.125\leq|\mathbf{s}|\leq 0.25\}$ the reference solid domain corresponding to the annulus at rest. The physical parameters are densities $\rho_{f}=1.0$ and $\rho_{s}=1.1$, equal fluid and solid viscosities $\nu=0.01$, and elasticity constant $\kappa=0.2$. We assume that at $t=0$ the fluid is at rest, hence $\mathbf{u}_{0}=0$, while the annulus is stretched and its initial deformation is given by $\mathbf{X}_{0}=\left({s_{1}}/{1.4}+0.5,1.4\,s_{2}+0.5\right)$. The internal elastic forces drive the annulus back to its resting position, thus generating the motion of the system, which is analyzed until $T=4$.
The mesh for the pressure unknown, $\mathcal{T}_{h}^{\Omega}$, is obtained by dividing $\Omega$ into $N^{2}$ squares then split into two triangles, with $h_{\Omega}=1/N$. Consequently, the velocity unknown is approximated on $\mathcal{T}^{\Omega}_{h/2}$, see (17). The grid $\mathcal{T}_{h}^{\mathcal{B}}$ is generated with Gmsh [27] in such a way that $h_{\mathcal{B}}\approx(4/3)h_{\Omega}$. We consider three configurations by setting $N=32\text{ (coarse)},\,64\text{ (medium)},\,128\text{ (fine)}$. For the time step, we choose $\Delta t=0.1,\,0.01$. For the coupling, we use $\mathbf{c}=\mathbf{c}_{0}$ computed exactly.
In Figure 14, some snapshots report the velocity contour lines, together with the position of the structure (top line) and the pressure profiles (bottom line). We plot the system configuration at time $t=0.5$, $t=1$ and $t=4$: this choice is due to the fact that the evolution of the immersed structure is faster at the beginning, while after $t=2$, it almost reaches the resting position. Moreover, since the pressure is discontinuous along the interface and it is discretized by continuous piecewise linear elements, the Gibbs phenomenon appears. Notice that the value of $p_{h}$ ranges in (0.05,0.05)(-0.05,0.05) at $t=0.5$, then it decreases to values in $(0,0.01)$ at $t=4$.
In order to assess the unconditional stability with respect to the time step, we compute the system energy
$\mathcal{E}(\mathbf{u}_{h}^{n},\mathbf{X}_{h}^{n})=\frac{\rho_{f}}{2}\|\mathbf{u}_{h}^{n}\|^{2}_{0,\Omega}+\frac{\delta\rho}{2}\left\|\frac{\mathbf{X}_{h}^{n}-\mathbf{X}_{h}^{n-1}}{\Delta t}\right\|_{0,\mathcal{B}}^{2}+\frac{\kappa}{2}|\mathbf{X}_{h}^{n}|_{1,\mathcal{B}}^{2}.$
Figure 15 displays the behavior of the energy ratio $\mathcal{E}(\mathbf{u}_{h}^{n},\mathbf{X}_{h}^{n})/\mathcal{E}(\mathbf{u}_{h}^{0},\mathbf{X}_{h}^{0})$, computed by setting $\mathbf{X}_{h}^{-1}=0$. We observe that the energy is decreasing in the interval of interest and again that the evolution is faster for $t\in(0,2)$. After that, the resting position is almost reached. Essentially, the energy behaves independently of the considered meshes and time steps. It seems that the possible presence of small cut cells does not have any effect on the energy. In order to check if the computation of the coupling term actually produces small cut cells, we follow the evolution of a randomly chosen solid element. At each time step, the picked element is subdivided into several (polygonal) cells taking into account the intersections with the underlying velocity mesh. We evaluate the area of these sub-cells and we mark it in the semi-log scatter plots in Figure 16. We see that, for $\Delta t=0.1$, cells with area of the order $^{-9}1010$$^{-10}$ appear for all meshes, while for $\Delta t=0.01$, the smallest cells can have an area ranging from $10^{-12}$ to $10^{-15}$, depending on the mesh refinement.
Figure 14. Snapshots of the annulus evolution (coarse test). Streamlines of $|\mathbf{u}_{h}|$ and structure position are reported on the top line, while the pressure field on the bottom line.
Figure 15. Energy ratio $\mathcal{E}(\mathbf{u}_{h}^{n},\mathbf{X}_{h}^{n})/\mathcal{E}(\mathbf{u}_{h}^{0},\mathbf{X}_{h}^{0})$ of the stretched annulus: evolution in time.
Figure 16. Area of the cut cells resulting from a fixed solid element. For each test, the solid element is randomly chosen and its evolution in time is studied.